How many ? and ? bonds are in this molecule?


How many sigma and pi bonds are in this molecule?
Answer #1
Concepts and reason

This problem is based on the concept of sigma and pi bonds present in compounds.

A chemical compound contains various bonds between the elemental atoms which involve sigma (σ)\left( \sigma \right) bonds as well as pi (Π)\left( \Pi \right) bonds. Sigma bonds are stronger than pi bonds as they are generated by head to head overlapping of orbitals whereas pi bonds are the result of side-ways overlapping of orbitals.

Fundamentals

An unsaturated compound contains double bonds and triple bonds. In case of double, one bond is sigma and the other is pi bond whereas in case of triple bonds, one bond is sigma and the other two bonds are pi bonds.

The structure of the compound is given below:

=ċ-
ċ-CSC-H

The structure of the compound is shown below:

Ans:

There are 9 sigma bonds (σ)\left( \sigma \right) and 4 pi (Π)\left( \Pi \right) bonds in the molecule.